OpenStudy (anonymous):

Find the most general f. Use C for the constant of the first anti-derivative and D for the constant of the second anti-derivative. f ''(x) = 4x + sin x

myininaya (myininaya):

\[\int\limits_{}^{}f''(x) dx=f'(x)+C\] \[\int\limits_{}^{}(f'(x)+C)dx=f(x)+Cx+D\] where C and D are constants so to find f you need to find f' first

OpenStudy (anonymous):

thank you for the procedure, myininaya. Do you have an answer?

myininaya (myininaya):

\[f'(x)=4*\frac{x^{1+1}}{1+1}-cosx+C\] \[f'(x)=2x^2-cosx+C\] \[f(x)=2*\frac{x^{2+1}}{2+1}-sinx+Cx+D\]

OpenStudy (anonymous):

Er...is it supposed to be "x^2+1" or did you intentionally write the numerator with "2+1" as the exponent for x?

myininaya (myininaya):

yes 2+1 is the exponent

myininaya (myininaya):

\[f(x)=2*\frac{x^3}{3}-sinx+Cx+D\] i was just showing you what rule i was using to integrate(showing you my work)

myininaya (myininaya):

hipster do you got it?

OpenStudy (anonymous):

Yep I got it. Thanks :)